With the wind, an airplane travels 2200 miles in 4 hours. Against the wind, it takes 5 hours to travel 2350 miles. Find the rate of the plane in still air.
Answer by Paul(988)

Let the rate of the wind be y
Let the rate of the plane be x
Equation 1:
4x+4y=2200
y=550-x (subsittuion)
Equation 2:
5x-5y=2350
Apply subsitution:
5x-5(550-x)=2350
10x=5100
x=510
y=550-510
y=40
Hence, the speed of the plane is 510mph and the speed of the wind is 40mph.
Paul.